# LETS-C: Leveraging Language Embedding for Time Series Classification

### v1

@@ Appendix C Model Performance
Figure 3: Model comparison based on classification accuracy, averaged across all datasets listed in Table 6. For detailed results, refer to Table 1.
Appendix D Computational Cost Analysis
Table 7 discusses the training and inference times for our model as compared to the SOTA OneFitsAll (Zhou et al., 2024) across the 10 benchmark datasets.
Table 7: Comparison of training and inference times (in seconds) per batch for OneFitsAll (Zhou et al., 2024) versus our model across the 10 benchmark datasets. The Delta (%) is calculated using the formula 100×Cost with Our ModelCost with $100\times\frac{\text{Cost with Our Model}}{\text{Cost with OneFitsAll}}$, quantifying the computational efficiency of our model relative to OneFitsAll.
@@ Appendix D Computational Cost Analysis
Time
Training time ($s$) LETS-C (Ours) 0.01 0.01 0.009 0.01 0.01 0.11 0.004 0.01 0.01 0.01 0.02
OneFitsAll 0.35 0.23 0.14 0.07 0.02 0.28 0.11 0.13 0.97 0.07 0.24
Delta (%) 3.17 8.02 6.71 16.18 42.85 40.49 4.30 8.40 1.21 15.30 14.66
Inference time ($s$) LETS-C (Ours) 0.01 0.05 0.01 0.01 0.01 0.10 0.004 0.01 0.02 0.01 0.02
OneFitsAll 0.16 0.10 0.06 0.04 0.02 0.15 0.06 0.08 0.15 0.04 0.09
Delta (%) 10.58 50.50 14.75 31.15 68.41 66.10 7.02 19.59 18.00 30.55 31.67
We observe that our approach significantly reduces the total training time to just 14.66% and the inference time to 31.67% of those required by OneFitsAll.
Appendix E Assessing Text Embeddings with Cosine Similarity
@@
Appendix F Numerical Precision for Tokenization
@@
Studies in the NLP domain have shown that longer inputs do not perform well with language models Press et al. (2020); Levy et al. (2024). Our empirical analysis supports this claim, revealing a decrease in classification accuracy with increased numerical precision when computing language embeddings. The average accuracy starts at 72.8% with precision 1 and declines to 69% at precision 6. Additionally, the percentage of AvgWins is highest at precision 1. As numerical precision increases, so does the length of the time series and the input to the language embedding model. Note that the maximum token length for text-embedding-3-large embeddings is 8191, and thus keeping precision of 1 ensures that context length doesn’t exceed the maximum permissible token length. This issue is especially problematic for datasets with longer time series, such as the EthanolConcentration dataset, which includes 1751 time steps per sample. This finding led us to opt for precision 1 in our study.
@@
Appendix G Embedding Type
Table 9 provides a summary of selected embedding models used in the study.
@@ Appendix G Embedding Type
Appendix H Trade-offs: Model Accuracy vs. Parameter Complexity
Table 10 illustrates the trade-off between model accuracy and the complexity of training parameters in our model, which utilizes both embeddings and time series data as inputs to a lightweight framework.
@@
This trade-off between model accuracy and parameter complexity is data-dependent. However, we generally observe a trend where a reduction in parameters leads to only a slight decrease in accuracy. Next, let’s take a closer look at three datasets: Heartbeat, PEMS-SF, and Spoken Arabic Digits, to understand how these trade-offs manifest in different contexts.

### v2

@@ I.2 Optimal Hyperparameters
Figure 3: Model comparison based on classification accuracy, averaged across all datasets listed in Table 4. For detailed results, refer to Table 1 in the main paper.
Appendix K Monetary Costs Analysis
Our approach incurs monetary costs when utilizing OpenAI’s text-embedding-3-large model OpenAI (2024). To compute the cost of generating text embeddings for our datasets, we followed the pricing structure of \$0.130 per million tokens. Each time series sample, including all time steps, was embedded separately for each dimension. To estimate the number of tokens per sample, we considered the numerical format of the time series values, where each time step consists of a three-digit number. These digits are separated by spaces, and consecutive time steps are delimited by commas, resulting in an estimated six tokens per time step. The total token count for each dataset was then computed as:
@@
[…]
where the estimated tokens per series were derived from the series length multiplied by six. The dataset characteristics, including train and test sizes, number of dimensions, and series lengths, are detailed in Table 4. Using this approach, the total number of tokens across all datasets was approximately 1.05 billion, leading to an estimated embedding cost of \$137.07. It is important to note that API costs may fluctuate over time, leading to potentially different or outdated values. Furthermore, cost remains a common limitation for all LLM-based approaches that rely on closed-source paid models.
Appendix L Assessing Text Embeddings with Cosine Similarity
@@
Appendix M Numerical Precision for Tokenization
@@
Studies in the NLP domain have shown that longer inputs do not perform well with language models Press et al. (2020); Levy et al. (2024). Our empirical analysis supports this claim, revealing a decrease in classification accuracy with increased numerical precision when computing text embeddings. The average accuracy starts at 72.8% with precision 1 and declines to 69% at precision 6. Additionally, the percentage of AvgWins is highest at precision 1. As numerical precision increases, so does the length of the time series and the input to the text embedding model. Note that the maximum token length for text-embedding-3-large embeddings is 8191, and thus keeping precision of 1 ensures that context length doesn’t exceed the maximum permissible token length. This issue is especially problematic for datasets with longer time series, such as the EthanolConcentration dataset, which includes 1751 time steps per sample. This finding led us to opt for precision 1 in our study.
@@
Appendix N Generalization Across Various Text Embedding Models
To evaluate the generalization capabilities of our approach across different text embedding models beyond text-embedding-3-large OpenAI (2024), the performance of LETS-C was tested using three alternative embedding models: e5-mistral-7b-instruct Wang et al. (2024), gte-large-en-v1.5 Li et al. (2023), and nomic-embed-text-v1 Nussbaum et al. (2024). A summary of the embedding models utilized in this study is provided in Table 8.
@@
N.1 e5-mistral-7b-instruct
The e5-mistral-7b-instruct model Wang et al. (2024) is based on the pretrained Mistral-7b checkpoint Jiang et al. (2023) and was fine-tuned using the RankLLaMA training methodology Ma et al. (2024). It employed LoRA Hu et al. (2021) with a rank of 16 on a mixture of multilingual datasets, which included both synthetic data and a collection of 13 public datasets, yielding approximately 1.8 million examples after sampling. Proprietary LLMs, such as GPT-4, were prompted to generate diverse synthetic data with instructions in multiple languages. Leveraging the strong language understanding capabilities of the Mistral model, e5-mistral-7b-instruct achieved state-of-the-art results across nearly all task categories on the competitive MTEB benchmark. Techniques such as gradient checkpointing, mixed precision training, and DeepSpeed ZeRO-3 were applied to further reduce GPU memory requirements.
N.2 gte-large-en-v1.5
The gte-large-en-v1.5 model Li et al. (2023) is a general text embedding (GTE) model that utilizes contrastive learning on an open-source large-scale dataset comprising unsupervised text pairs extracted from various sources. To enhance the quality of the learned text representations, high-quality text pairs with human labels from multiple sources were employed for contrastive fine-tuning. The model utilizes a multi-stage contrastive learning approach and benefits from a diverse training data mixture, enabling it to achieve strong generalization performance for single-vector embeddings.
N.3 nomic-embed-text-v1
The nomic-embed-text-v1 model Nussbaum et al. (2024) is a fully reproducible, open-source English text embedding model with an 8192 context length that outperforms both OpenAI Ada-002 and OpenAI text-embedding-3-small on short and long-context tasks. To accommodate long sequence lengths, the model adapts the BERT architecture Devlin et al. (2019) with several optimizations: replacing absolute positional embeddings with rotary positional embeddings Su et al. (2024), using SwiGLU activation instead of GeLU Shazeer (2020), implementing Flash Attention Dao et al. (2022), setting Dropout to 0 Geiping and Goldstein (2023), and ensuring the vocabulary size is a multiple of 64 Portes et al. (2024); Shoeybi et al. (2019). These modifications result in a 137M parameter encoder.
Appendix O Trade-offs: Model Accuracy vs. Parameter Complexity
Table 9 illustrates the trade-off between model accuracy and the complexity of training parameters in our model, which utilizes both embeddings and time series data as inputs to a lightweight framework. To vary model size, we adjust the number of linear and convolution layers in the classification head, ranging from 1 to 5 layers each, creating model variants of different sizes.
@@
Across all datasets, we find that significant reductions in model parameters result in only a minimal loss of accuracy. This trade-off between model accuracy and parameter complexity is data-dependent. However, we generally observe a trend where a reduction in parameters leads to only a slight decrease in accuracy. Next, let’s take a closer look at three datasets: Heartbeat, PEMS-SF, and Spoken Arabic Digits, to understand how these trade-offs manifest in different contexts.
@@
Appendix P Ablation Study
Table 10 presents the results of this study, comparing the performance of our default approach, which adds embeddings to time series, to variants that exclude either embeddings or the time series.
[…]
LETS-C (Ours) 52.9 68.9 23.8 78 99.2 93.1 93.2 62.8 99.2 90.6 76.17 60%
   embedding only 38 59.4 20.1 80 99.2 92.5 88.7 63.9 99.2 88.4 72.94 40%
Appendix Q Alternative Methods for Fusing Time Series with Embeddings
We explored two additional methods for fusing time series and embeddings beyond simple addition. The first method involves a Fusion network that first processes embeddings and time series data through convolutional and dense layers in two separate branches, then merges the features from both branches into a final dense network. The second method employs Concatenation, where the time series and embeddings are concatenated and processed through a lightweight classification head. Despite cross-attention being another alternative for fusing different modalities, we didn’t include it in this study due to the computational complexity it adds to the model. Table 11 presents the classification accuracy and trainable model parameters for these variations.
Table 11: Comparison of classification accuracy (%) and trainable model parameters (millions) for alternative methods of fusing time series with embeddings. Higher AvgWins and accuracy averages indicate superior performance, while lower model parameter averages suggest greater computational efficiency. Red: Best performance.
[…]
Concatenation 43 65.1 22.5 79 98.9 93.1 93.9 58.9 99 88.8 74.22 30%
[…]
Fusion Network 0.19 0.35 0.33 0.28 0.16 5.54 0.37 0.27 0.23 0.04 0.78
Concatenation 0.42 0.009 0.26 0.17 0.11 0.28 0.23 0.39 0.09 0.46 0.24
We observe that the addition approach in the LETS-C architecture achieves the highest average classification accuracy (76.11%) compared to the fusion network (73.40%) and concatenation (74.22%). Notably, when compared to all baselines, concatenation emerges as the second-best method after addition, which achieves the highest accuracy. The fusion approach ranks fifth overall, with only OneFitsAll and TimesNet outperforming it, following addition and concatenation, in terms of average accuracy.